# Universal control of four singlet-triplet qubits

### v2

@@ Acknowledgements.
We thank C. Déprez for insightful discussions and kind help. We also thank other members of the Vandersypen group and Veldhorst group for their stimulating discussions. We acknowledge S. L. de Snoo’s help in software development and technical support by O. Benningshof, J. D. Mensingh, R. Schouten, R. Vermeulen, R. Birnholtz, E. Van der Wiel, and N. P. Alberts. This work was funded by an Advanced Grant from the European Research Council (ERC) under the European Union’s Horizon 2020 research. M.R.-R. acknowledges support from the Netherlands Organization of Scientific Research (NWO) under Veni grant VI.Veni.212.223.
@@
Data availibility The data reported in this paper are archived on a Zenodo data repository at https://doi.org/10.5281/zenodo.10431402

### v3

@@ Acknowledgements.
We thank C. Déprez, I. Fernandez De Fuentes, B. Undseth, Y. Matsumoto and X.X. Yang for insightful discussions and kind help. We also thank other members of the Vandersypen group and Veldhorst group for stimulating discussions. We acknowledge S. L. de Snoo’s help in software development and technical support by O. Benningshof, J. D. Mensingh, R. Schouten, R. Vermeulen, R. Birnholtz, E. Van der Wiel, and N. P. Alberts. This work was funded by an Advanced Grant from the European Research Council (ERC) under the European Union’s Horizon 2020 research (882848). C.-A.W. and M.V. acknowledge support by the European Union through an ERC Starting Grant QUIST (850641). M.R.-R. acknowledges support from the Netherlands Organization of Scientific Research (NWO) under Veni grant VI.Veni.212.223.
@@
Data and Code availability The data supporting this work and codes used for data processing and numerical simulation are archived on a Zenodo data repository at https://doi.org/10.5281/zenodo.12801188